# Entropy-Guided Reasoning Compression

### v1

@@ Abstract
Hourun Zhu   Yang Gao thanks: Corresponding author.   Wenlong Fei   Jiawei Li   Huashan Sun School of Computer Science and Technology Beijing Institute of Technology, Beijing, China {hrzhu, gyang, wenlongfei, jwli, hssun }@bit.edu.cn
@@ 3.1 Preliminary
With the introduction of a compression constraint, it indicates that compression narrows the output space and reduces entropy. Conversely, when accuracy-oriented rewards dominate, the model tends to explore longer reasoning paths to improve accuracy, which leads to an increase in entropy. This reveals an intrinsic opposition between accuracy optimization and reasoning compression—an entropy conflict that underlies instability in joint training.
@@ 4.1 Experimental Setups
We use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B as the base model and perform reinforcement learning training on the DeepScaleR-Preview deepscaler2025 dataset. The detailed training configurations for both models are provided in the appendix. We evaluate the trained models on six mathematical benchmarks, measuring both Pass@1 accuracy and response length. The benchmarks include: AIME24 aime24, AMC, MATH math500, MinervaMATH Minerva, Olympiad-Bench olympiadbench, GSM-8K gsm8k. To better assess the model’s reasoning efficiency, we introduce a key metric for evaluation from bingo.
@@ 4.3 Empirical Analysis of Entropy Conflict in Reasoning Compression
Exploration → Compression (after exploration) 40.0 6721 97.3 982 87.6 1992 86.8 714
@@
Figure 4: Comparison of reasoning length and correctness on 30 AIME24 samples. Each bar shows the number of generated tokens for a sample, with symbol indicating correctness and the number on top denoting the count of reasoning steps. The compressed model produces substantially shorter reasoning chains while maintaining comparable accuracy.
@@ 4.5 Case Study
This case study compares the reasoning behavior of the original DeepSeek model and our compressed model on 30 AIME24 problems. The original model often produces long chains with many intermediate steps, resulting in thousands of tokens per sample. In contrast, the compressed model generates much shorter chains across almost all samples. Despite the large reduction in token count and reasoning steps, correctness remains largely aligned between the two models. This shows that long reasoning is not always necessary for solving these problems and that our method preserves essential reasoning structure while removing redundant steps. The result highlights the effectiveness of compression in eliminating unnecessary expansion without undermining the model’s mathematical reasoning ability.

### v2

@@ Abstract
Hourun Zhu   Yang Gao thanks: Corresponding author.   Wenlong Fei   Jiawei Li   Huashan Sun School of Computer Science and Technology Beijing Institute of Technology, Beijing, China {zhuhr, gyang, wenlongfei, jwli, hssun }@bit.edu.cn
@@ 3.1 Preliminary
With the introduction of a compression constraint, it indicates that compression narrows the output space and reduces entropy. Conversely, when accuracy-oriented rewards dominate, the model tends to explore longer reasoning paths to improve accuracy, which leads to an increase in entropy. This reveals an intrinsic opposition between accuracy optimization and reasoning compression—an entropy conflict that underlies instability in joint training. Further discussion can be found in B.
@@ 4.1 Experimental Setups
We use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B as the base model and perform reinforcement learning training on the DeepScaleR-Preview deepscaler2025 dataset. The detailed training configurations for both models are provided in the appendix A. We evaluate the trained models on six mathematical benchmarks, measuring both Pass@1 accuracy and response length. The benchmarks include: AIME24 aime24, AMC, MATH math500, MinervaMATH Minerva, Olympiad-Bench olympiadbench, GSM-8K gsm8k. To better assess the model’s reasoning efficiency, we introduce a key metric for evaluation from bingo.
@@ 4.3 Empirical Analysis of Entropy Conflict in Reasoning Compression
Exploration → Compression (after exploration) 40.0 15127 97.5 2953 88.6 5135 87.9 2202
@@
(a) Token Distribution by Group
(b) Reasoning Steps Distribution by Group
Figure 4: Comparison of the original model and our compressed model on the Math500 benchmark, grouped by four correctness transitions: preserved (✓→✓), lost (✓→×), gained (×→✓), and failed (×→×). The boxplots show the distribution of (a) token counts and (b) reasoning steps for both models within each group.
@@ 4.5 Case Study
Our results on Math500 reveal that the proposed compression method not only shortens the generated reasoning traces in terms of token length, but also effectively removes redundant reasoning steps. As shown in Figure 4, the responses of the original model are already fairly concise on problems it answers correctly, indicating that it does not engage in excessive or unnecessary reasoning when it can handle the problem confidently. Our training further compresses these already-short traces while largely preserving correctness. However, this compression can occasionally remove useful intermediate reasoning, leading to a small portion of “✓→×” cases where correct answers become incorrect.
Interestingly, we also observe the opposite trend: a nontrivial number of samples fall into the “×→✓” group. After compression, the model becomes more accurate because eliminating noisy or misleading reasoning steps helps the model focus on the essential logical structure of the problem. This phenomenon suggests that redundant reasoning is not merely inefficient—it may actually harm accuracy. Further inspection C reveals that the compressed model has acquired a more efficient reasoning strategy. During the internal think phase, it no longer performs explicit calculations; instead, it produces only high-level procedural guidance. The concrete computations are carried out in the final output stage, where the model executes the necessary steps rapidly and with minimal redundancy. This behavioral shift indicates that the model has learned to separate planning from execution, leading to substantially more concise and purpose-driven reasoning traces.
@@ 5 Conclusion
Appendix A Implementation Details
We implement our training pipeline based on the Verl framework verl. During optimization, we use a batch size of 64 and update the model every 16 samples, with eight sampled responses per input to estimate the policy gradient. We adopt the AdamW optimizer with a learning rate of 1e-6. The entropy-guided training consists of two consecutive stages: a compression stage of 400 steps followed by an exploration stage of 400 steps.
Figure 5: Case study on AIME24.
Appendix B Discussion
Although our work focuses on the compression setting, the entropy-conflict analysis provides a more general perspective for understanding multi-objective optimization in large-scale model training. As discussed in §3.1, entropy conflict reflects a structural limitation: when a model is optimized toward a subset of objectives, it continues to preserve gradients associated with the original broader objective set, preventing full convergence to the subspace even when the new objective dominates.
This insight suggests broader implications for RL-based LLM training pipelines. Many hybrid reinforcement learning objectives implicitly form nested objective sets—some goals are strict subsets of others. In such cases, jointly optimizing mixed objectives without considering this hierarchy may introduce gradient interference and slow specialization. A more efficient strategy is to explicitly exploit the inclusion structure: first train on the sub-objective set to let the model acquire the core behavioral rules, and then train on the superset to encourage free exploration under these learned constraints. Conceptually, the subset stage establishes the “rules of the game,” and the superset stage enlarges the space for exploration while ensuring the model remains anchored to the desired behaviors. This staged perspective offers a principled approach for improving controllability and sample efficiency in real-world model development.
Appendix C Case study
In this example(Figure 5), the original model performs substantial computation inside the think phase: it repeatedly carries out algebraic manipulations, reevaluates earlier steps, and even restarts partial derivations before producing its final answer. After the think block ends, the model effectively recomputes the solution a second time in the visible output, resulting in severe redundancy and inflated reasoning length. In contrast, our compressed model adopts a fundamentally different behavior pattern. During the think phase it provides only high-level guidance—structuring the solution path without executing detailed calculations. All concrete computation is deferred to the final answer stage. This separation removes unnecessary internal loops, shortens the reasoning chain by eliminating duplicated work, and produces a more efficient and interpretable solution trajectory.